# SiamSeg: Self-Training with Contrastive Learning for Unsupervised Domain Adaptation in Remote Sensing

### v1

@@ Abstract
Semantic segmentation of remote sensing (RS) images is a challenging task that holds significant potential across various applications. Deep learning, particularly supervised learning utilizing large-scale labeled datasets, has markedly advanced this field. However, acquiring high-quality labeled data is prohibitively expensive and time-consuming. Additionally, variations in ground sampling distance (GSD), remote sensing imaging equipment, and geographic diversity contribute to a pervasive domain shift between datasets. This domain shift poses substantial challenges to supervised learning methods that depend on annotated source domain data, resulting in poor performance when processing cross-domain RS images. It is well-established that the domain shift problem significantly undermines a model’s generalization ability in the target domain. To mitigate this issue, we explore the unsupervised domain adaptation (UDA) approach. This method aims to concurrently learn from unlabeled target domain data while training on labeled source domain data. Unlike traditional adversarial learning methods, recent advancements in self-supervised learning through pseudo-label generation for the target domain have shown promising results. Various strategies have been proposed to address the low-quality pseudo-labeling issues stemming from domain bias. Among these, the technique of integrating source and target domain images along with their corresponding true and pseudo labels for self-supervised training is emerging as an effective solution to combat domain discrepancies. Despite the progress of pseudo-labeling methods in computer vision, their application to RS image semantic segmentation tasks remains underexplored. Consequently, this paper investigates a self-supervised training (ST) approach based on pseudo-label generation. We note that existing pseudo-label generation methods typically select only high-confidence pixel points as pseudo-labels, which diminishes the effectiveness of supervised signals in low-confidence regions. Furthermore, self-supervision methods that rely on pseudo-labeling are limited in their ability to learn semantic information about the target domain due to the absence of ground truth labels (GTs) as supervised data. Although some methods attempt to leverage the overall confidence level as pseudo-labeling weights, noise (i.e., incorrect pseudo-label assignments) can still arise during pseudo-label generation, further compromising the network’s ability to learn semantic information from the target domain. With the emergence of contrastive learning in computer vision, this approach has proven effective in capturing semantic information from images. To tackle the challenge of insufficient model representation in UDA due to the lack of labels for target domain data, we propose integrating a contrastive learning network into UDA, which provides additional supervised information to enhance the model’s visual recognition capabilities in the target domain. Specifically, the model learns semantic information from the same target domain image by maximizing similarity between its augmented versions, thereby improving the model’s representational capacity. Extensive experiments conducted on several well-known remote sensing datasets (Potsdam, Vaihingen, and LoveDA) demonstrate that the proposed SimSeg method significantly outperforms existing approaches, achieving state-of-the-art performance levels. Visualization results and quantitative analyses further validate that SimSeg excels in learning within the target domain. The code is publicly available at https://github.com/woldier/SiamSeg .

### v2

# SiamSeg: Self-Training with Contrastive Learning for Unsupervised Domain Adaptation Semantic Segmentation in Remote Sensing

@@ Abstract
Semantic segmentation of remote sensing (RS) images is a challenging yet crucial task. While deep learning, particularly supervised learning with large-scale labeled datasets, has significantly advanced this field, acquiring high-quality labeled data is expensive and time-consuming. Additionally, variations in ground sampling distance, imaging equipment, and geographic differences cause domain shifts between datasets, which limit model performance across domains. Unsupervised domain adaptation (UDA) offers a solution by enabling models to learn from unlabeled target domain data while training on labeled source domain data. Recent self-supervised learning approaches using pseudo-label generation have shown promise in addressing domain discrepancies. Combining source and target images with their true and pseudo labels has proven effective in reducing domain bias. However, the use of pseudo-labeling for RS image segmentation is underexplored. Existing methods often rely on high-confidence pixel points as pseudo-labels, reducing supervision in low-confidence areas. Noise in pseudo-labels further weakens the model’s ability to learn target domain semantics. While some methods assign confidence weights, noisy pseudo-labels remain an issue. To address these limitations, we propose integrating contrastive learning into UDA, enhancing the model’s capacity to capture semantic information by maximizing the similarity between augmented views of the same image. This provides additional supervision to improve performance in the target domain. Extensive experiments on key RS datasets (Potsdam, Vaihingen, LoveDA) demonstrate that our SimSeg method outperforms existing approaches, achieving state-of-the-art results. Visualization and quantitative analyses confirm its superior ability to learn from the target domain. The code is available at https://github.com/woldier/SiamSeg.